If Leah is 6 years older than her sister, Sue, and John is 5 years older than Leah, and the rotal of their ages is 41.
6+5+6=41

Answer by rapaljer(4551)

Let x = Sue's age
x+6 = Leah's age
x+6 + 5 = John's age

Sum of their ages=
x + (x+6) + (x+6+5) = 41
3x +17 = 41

Subtract 17 from each side:
3x + 17 - 17 = 41 - 17
3x = 24

Divide both sides by 3:
x= 8 Sue's age
x+6 = 14 Leah's age
x + 11 = 19 = John's age

Check: Sum of their ages = 8+14+19 = 41

R^2 at SCC